Find the value of x for which (2+x)⁴=16 + 16x + x⁴
Could I be shown this step by step please ''.

Answer by stanbon(75887)

Find the value of x for which (2+x)⁴=16 + 16x + x⁴
--------
4C4*2^4 + 4C3*2^3x + 4C2*2^2*x^2 + 4C1*2*x^3 + 4C0*x^4 = 16 + 16x + x^4
----------------------------------------------------
16 + 32x + 24x^2 + 8x^3 + x^4 = 16 +16x+x^4
-----------------------------------------------
16x + 24x^2 + 8x^3 = 0
-------------------------------
8x(2 + 3x + x^2) = 0
8x(x+2)(x+1) = 0
x = 0 or x = -1 or x = -2
---------------------------
Cheers,
Stan H.
-----------